# Automated threshold selection and associated inference uncertainty for univariate extremes

### v3

@@ Acknowledgements
Case 4 $\bm{0.108}$ $\bm{-0.077}$ 0.006 0.111 -0.083 $\bm{0.005}$
Table S.6: RMSE, bias and variance of threshold estimates for Cases 1-4 for the EQD method found using the different calibration data in the metric $d_{b}(u)$: the proposed EQD with $Q(p,\bm{x}^{b}_{u})$ and the adjusted version with $Q(p,\bm{x}_{u})$.
@@
We provide results that expand on those in Section 6 of the main text. Section S:6.2 provides bias-variance decompositions for the RMSEs values given in Section 6 and presents additional results for the Danielsson et al., (2001, 2019) methods, evaluating threshold estimation for Cases 1-4 and quantile estimation for Gaussian data. In Section S:6.3, we provide further threshold estimation results for additional cases, which have alternative parameters and sample sizes to those of Cases 1-4.
S:6.2 Details results for the case studies
@@
For Cases 1-4, Table S.7 shows the RMSE, bias and variance of the thresholds selected by the EQD, Wadsworth, (2016) and Northrop et al., (2017) methods. The RMSE is also reported in Table 2 of the main paper, but is repeated here for completeness. The EQD method has the smallest RMSE and the least variable estimates in all cases, and is the least biased method in 3 of the 4 cases. Table S.8, presents equivalent results for the Danielsson et al., (2001, 2019) methods. Both methods show considerably larger RMSEs than the methods presented in Table S.7, due to the large positive biases of these methods across all cases. In particular, for Cases 3 and 4 the Danielsson et al., (2019) method has the smallest variance of all the methods but its larger bias leads to RMSE values much larger than those of the EQD and other methods in Table S.7.
@@
Case 4 $\bm{0.526}$0.515 $\bm{0.012}$ 0.627 -0.407 0.230 0.543 $\bm{-0.222}$ 0.246
Table S.7: RMSE, bias and variance of the threshold estimates for the EQD, Wadsworth and Northrop methods for Cases 1-4.
@@
Table S.8: RMSE, bias and variance of the threshold estimates for the Danielsson et al., (2001, 2019) methods used for Cases 1-4.
@@
Tables S.9 & S.10 present the bias and variance of the quantile estimates for the EQD, Wadsworth, (2016) and Northrop et al., (2017) methods applied to Cases 1-4. These are shown for the $(1-p_{j})$-quantile estimates where $p_{j}=1/(10^{j}n)$ for $j=0,1,2$ and $n$ denotes the length of the simulated dataset. As mentioned in Section 6 of the main paper, we use exceedance probabilities of this form because we have simulated samples of different sizes and want to make extrapolation equally difficult in each case. These bias and variance components correspond to the RMSE values presented in Table 3 in Section 6 of the main paper, where the EQD method achieves the lowest RMSEs in all cases and quantiles. Table S.10 shows that these lower RMSE values derive mainly from the variance component; for all $j$ and in all cases the EQD method shows the least variability in quantile estimates, with the differences between the methods becoming more evident for higher $j$. The smallest absolute bias values in Table S.9 vary between each of the methods but the EQD incurs the least bias in the majority of cases and quantiles. In particular, the EQD achieves the smallest absolute bias values for all $j$ in Cases 2 and 3.
EQD Wadsworth5 Northrop EQD Wadsworth5 Northrop
@@
EQD Wadsworth5 Northrop EQD Wadsworth5 Northrop
@@
We next consider the case of Gaussian data, using 500 simulated datasets of size $n=2000$ and $20000$. Tables S.12 and S.13 present the bias, variance and RMSE for the estimation of the $(1-p_{j})$-quantiles (where $p_{j}=1/(10^{j}n)$ and $j=0,1,2$), based on the thresholds selected by the EQD, Wadsworth, (2016) and Northrop et al., (2017) methods. These RMSE values are detailed in Table 5 of the main paper.
@@
Table S.12: RMSE, bias and variance of estimated $(1-p_{j})$-quantiles from a Gaussian distribution with sample size of $2,000$: for EQD, Wadsworth and Northrop methods (top) and Danielsson et al., (2001, 2019) (bottom). Results are based on 500 replicated samples.
@@
Table S.13: RMSE, bias and variance of estimated $(1-p_{j})$-quantiles from a Gaussian distribution with sample size of $20,000$. Results are based on 500 replicated samples.
@@
There is no true GPD threshold for the Gaussian scenario but the quantile estimates discussed in Tables S.12 and S.13 require a preceding step of selecting a suitable threshold above which the GPD approximation is adequate. In Table S.14, we provide information about the selected thresholds for the Gaussian data, presenting the 2.5%, 50%, 97.5% values of the sampling distribution of the threshold estimates (presented as a quantile of the Gaussian distribution for each method). The results show that the Wadsworth, (2016) method tends to estimate the threshold lowest, followed by the EQD method, and then the Northrop et al., (2017) method which tends to estimate the highest threshold values, with this finding being consistent across the sampling distribution quantiles. It is interesting to see that, even with a very large sample size, almost always thresholds are estimated to be below the 95% quantile (the maximum of the candidate thresholds) of the Gaussian distribution. This is surprising given its widely known slow convergence issues.
@@
EQD Wadsworth6 Northrop EQD Wadsworth Northrop
@@
In all of the cases, results presented here, for each of the methods, are based on 500 replicated samples and we take the set of candidate thresholds as the sample quantiles at levels 0%, 5% …, 95%, as in Section 6 of the main paper. Specifically, for Case 8, we considered an additional finer grid of candidate thresholds as the sample quantiles at levels 0%, 0.5%, …, 95%. For Case 0, we initially used a set of candidate thresholds which contains values lower than the minimum of the sample, but Wadsworth, (2016) and Northrop et al., (2017) methods had major problems, and so we omit those results. However, this restriction to consider only candidate thresholds which are sample quantiles automatically positively biases threshold estimates in Case 0. In any simulation study, it is reasonable to consider candidate thresholds above and below the true threshold, so the fact that the EQD method continues to work well for candidate thresholds below the sample quantiles is a particularly pleasing feature, even if not illustrated here.
@@
Tables S.19 and S.20 expanding on the results given in Tables 4 and 6 of the main paper, providing coverage probabilities and average CI width ratios using Algorithms 1 and 2 for 500 samples derived from Case 4 and the Gaussian distribution. These results allow us to conclude that the additional threshold uncertainty, captured by Algorithm 2, leads to a significant improvement in the calibration of CIs and that it is vital to include this uncertainty in any inference for extreme quantiles.
@@
Table S.19 shows that for all confidence levels and exceedance probabilities, Algorithm 1, which includes uncertainty only from the GPD parameter estimation, substantially under-estimates the uncertainty in quantile estimates. This leads to CIs which do not cover the true quantiles to the nominal levels of confidence. In contrast, the inclusion of the additional threshold uncertainty in Algorithm 2 leads to significant increases in the coverage of true quantiles with coverage probabilities at all quantile levels lying very close to the nominal confidence level, especially at the 95% level. From a practical perspective, it is reassuring that this improvement in coverage is achieved with only 40-66% average increases in the width of the CIs.
@@
Table S.20 shows that, for Gaussian variables across all quantiles, Algorithm 1 and 2 are less successful in the coverage of the true quantiles than in Case 4. Specifically, there is a significant under-estimation of the estimated uncertainty necessary to provide coverage probabilities near to the nominal confidence level, and the actual coverage is decreasing with the level of extrapolation required. This is not too surprising as it is well-established that Gaussian variables exhibit quite slow convergence to an extreme value limit. However, what Table S.20 shows is that the additional threshold uncertainty in Algorithm 2 leads to a substantial improvement in actual coverage across all quantiles and for all nominal confidence levels. This improvement is achieved with the CI widths on average being extended by 44-70%. In particular, for the $p=1/n,1/3n,1/5n$, which are typical levels of extrapolation from a sample in practical contexts, Algorithm 2 achieves a workable performance, with coverage reasonably close to the nominal level.

### v4

@@ Acknowledgements
Case 4 $\bm{0.526}$ $\bm{-0.515}$ 0.012 0.545 $-0.535$ $\bm{0.011}$
Table S.6: RMSE, bias and variance of threshold estimates for Cases 1-4 for the EQD method found using the different calibration data in the metric $d_{b}(u)$: the proposed EQD with $Q(p_{j};\bm{x}^{b}_{u},\bm{q})$ and the adjusted version with $Q(p_{j};\bm{x}_{u},\bm{q})$.
S:5.5 Investigating the effect of bootstrapping
In this section, we provide further simulation experiments to investigate the effect of the bootstrapping component of the EQD method. We utilise a variant of the EQD method with no bootstrapping and compare the results against the original EQD method across Cases 0-4 and the Gaussian case (with $n=2000$), with each case based on 500 replicated samples (Case 0 is outlined in Section S:6.3, while Cases 1-4 and the Gaussian case are outlined in Section 6 of the main paper.).
Table S.7 provides results for the RMSE, bias and variance of threshold choices for the EQD method with and without bootstrapping, for Cases 0-4. For Cases 0-3, removing the bootstraps and only evaluating the metric on the original sample leads to RMSEs of threshold choice almost twice as large than for the original EQD method. In each of these cases, there is an increase in positive bias and the variance increases by a factor of at least 4. Thus, removing the bootstrapping component leads to higher and more variable threshold choices for these cases as would be expected. For Case 4, the removal of the bootstrapping component actually leads to a slight decrease in RMSE, in contrast to the previous cases. This decrease stems from a reduction in the negative bias component which is as a result of threshold choices being slightly higher than the original method. This reduction in bias comes at a cost of greater variability in selected thresholds. This greater variability is to be expected by not averaging over bootstrap samples.
Case 0 Case 1 Case 2
RMSE Bias Variance RMSE Bias Variance RMSE Bias Variance
Original 0.042 0.019 0.001 0.048 0.034 0.001 0.060 0.031 0.003
No bootstraps 0.095 0.039 0.008 0.090 0.054 0.005 0.122 0.063 0.011
Case 3 Case 4
RMSE Bias Variance RMSE Bias Variance
Original 0.060 0.042 0.002 0.526 $-0.515$ 0.012
No bootstraps 0.138 0.080 0.013 0.473 $-0.441$ 0.030
Table S.7: RMSE, bias and variance of threshold choice for Cases 0-4 for EQD only evaluated on original sample (i.e., no bootstrapping).
Table S.8 shows the RMSE of quantile estimation following threshold selection using the original EQD method and the variant with no bootstrapping for the Gaussian case. There is a slight improvement in performance in terms of RMSE from including the bootstrapping component across all quantiles but results are very similar across both methods.
$j$ Original No bootstraps
0 0.214 0.217
1 0.430 0.435
2 0.703 0.715
Table S.8: RMSEs of estimated $(1-p_{j,n})$-quantiles where $p_{j,n}=1/(10^{j}n)$, for $j=0,1,2$, for the Gaussian case.
Overall, across the studied cases, the addition of bootstrapping to the EQD method leads to a systematic reduction in the variance and lower threshold choices, typically resulting in reduced RMSE, for the selected thresholds and the subsequent quantile estimates. While the mean absolute deviation is a robust and effective metric for threshold selection, the bootstrapping component provides further stability in the threshold choices and allows us to account for the increasing uncertainty in parameter estimates as the threshold increases. These investigations support our choice to use bootstrapping as a key component of the EQD method.
@@
We provide results that expand on those in Section 6 of the main text. Section S:6.2 provides bias-variance decompositions for the RMSEs values given in Section 6 and presents additional results for the Danielsson et al., (2001, 2019) methods, evaluating threshold estimation for Cases 1-4 and quantile estimation for Gaussian data. In Section S:6.3, we provide further threshold estimation results for additional cases, which have alternative parameters and sample sizes to those of Cases 1-4. Section S:6.4 presents additional results to assess the sensitivity of the methods to the choice of candidate threshold grids using Cases 1 and 4 with candidate grids defined above the mode of the distribution and for the Gaussian case with candidate grids spanning the entire range of the sample.
S:6.2 Detailed results for the case studies
@@
For Cases 1-4, Table S.9 shows the RMSE, bias and variance of the thresholds selected by the EQD, Wadsworth, (2016) and Northrop et al., (2017) methods. The RMSE is also reported in Table 2 of the main paper, but is repeated here for completeness. The EQD method has the smallest RMSE and the least variable estimates in all cases, and is the least biased method Cases 1-3. Table S.9 also presents equivalent results for the Danielsson et al., (2001, 2019) methods. Both methods show considerably larger RMSEs than the EQD, Wadsworth, (2016) and Northrop et al., (2017) methods, due to the large positive biases of these methods across all cases. In particular, for Cases 3 and 4, the Danielsson et al., (2019) method has the smallest variance of all the methods but its larger bias leads to RMSE values much larger than those of the EQD and other methods in Table S.9.
@@
Case 4 $\bm{0.526}$ $-0.515$ $\bm{0.012}$ 0.627 $-0.407$ 0.230 0.543 $\bm{-0.222}$ 0.246
@@
Table S.9: RMSE, bias and variance of the threshold estimates for Cases 1-4: for EQD, Wadsworth and Northrop methods (top) and Danielsson et al., (2001, 2019) (bottom). Results are based on 500 replicated samples.
@@
Tables S.10 & S.11 present the bias and variance of the quantile estimates for the EQD, Wadsworth, (2016) and Northrop et al., (2017) methods applied to Cases 1-4. These are shown for the $(1-p_{j,n})$-quantile estimates where $p_{j,n}=1/(10^{j}n)$ for $j=0,1,2$ and $n$ denotes the length of the simulated dataset. As mentioned in Section 6 of the main paper, we use exceedance probabilities of this form because we have simulated samples of different sizes and want to make extrapolation equally difficult in each case. These bias and variance components correspond to the RMSE values presented in Table 3 in Section 6 of the main paper, where the EQD method achieves the lowest RMSEs in all cases and quantiles. Table S.11 shows that these lower RMSE values derive mainly from the variance component; for all $j$ and in all cases the EQD method shows the least variability in quantile estimates, with the differences between the methods becoming more evident for higher $j$. The smallest absolute bias values in Table S.10 vary between each of the methods but the EQD incurs the least bias in the majority of cases and quantiles. In particular, the EQD achieves the smallest absolute bias values for all $j$ in Cases 2 and 3.
EQD Wadsworth4 Northrop EQD Wadsworth4 Northrop
@@
EQD Wadsworth4 Northrop EQD Wadsworth4 Northrop
@@
We next consider the case of Gaussian data, using 500 simulated datasets of size $n=2000$ and $20000$. Tables S.13 and S.14 present the bias, variance and RMSE for the estimation of the $(1-p_{j,n})$-quantiles (where $p_{j,n}=1/(10^{j}n)$ and $j=0,1,2$), based on the thresholds selected by the EQD, Wadsworth, (2016) and Northrop et al., (2017) methods. These RMSE values are detailed in Table 5 of the main paper.
@@
Table S.13: RMSE, bias and variance of estimated quantiles from a Gaussian distribution with sample size of $2000$: for EQD, Wadsworth and Northrop methods (top) and Danielsson et al., (2001, 2019) (bottom). Results are based on 500 replicated samples.
@@
Table S.14: RMSE, bias and variance of estimated quantiles from a Gaussian distribution with sample size of $20000$. Results are based on 500 replicated samples.
@@
There is no true GPD threshold for the Gaussian scenario but the quantile estimates discussed in Tables S.13 and S.14 require a preceding step of selecting a suitable threshold above which the GPD approximation is adequate. In Table S.15, we provide information about the selected thresholds for the Gaussian data, presenting the 2.5%, 50%, 97.5% values of the sampling distribution of the threshold estimates (presented as a quantile of the Gaussian distribution for each method). The results show that the Wadsworth, (2016) method tends to estimate the threshold lowest, followed by the EQD method, and then the Northrop et al., (2017) method which tends to estimate the highest threshold values, with this finding being consistent across the sampling distribution quantiles. It is interesting to see that, even with a very large sample size, almost always thresholds are estimated to be below the 95% quantile (the maximum of the candidate thresholds) of the Gaussian distribution. This is surprising given its widely known slow convergence issues.
@@
EQD Wadsworth5 Northrop EQD Wadsworth Northrop
@@
In all of the cases, results presented here, for each of the methods, are based on 500 replicated samples and we take the set of candidate thresholds as the sample quantiles at levels 0%, 5% …, 95%, as in Section 6 of the main paper. Specifically, for Case 8, we considered an additional finer grid of candidate thresholds as the sample quantiles at levels 0%, 0.5%, …, 95%. For Case 0, we initially used a set of candidate thresholds which contains values lower than the minimum of the sample, but Wadsworth, (2016) and Northrop et al., (2017) methods had major problems, and so we omit those results. However, this restriction to consider only candidate thresholds which are sample quantiles automatically positively biases threshold estimates in Case 0. In any simulation study, it is reasonable to consider candidate thresholds above and below the true threshold, so the fact that the EQD method continues to work well for candidate thresholds below the sample quantiles is a particularly pleasing feature, even if not illustrated here. In Section S:6.4, we explore the effect of candidate thresholds which lie below the mode of the data for Cases 1, 4 and the Gaussian case for the EQD, Wadsworth and Northrop methods.
@@
S:6.4 Sensitivity to the choice of candidate threshold grids
In this section, we provide further simulation experiments where we choose the candidate threshold grids more in line with general extreme value analyses (i.e., above the mode) and explore the sensitivity of methods to candidate thresholds which lie below the mode of the data. Since the GPD density is monotonically decreasing for realistic values of $\xi$ (i.e., when $\xi>-1$), when utilising a threshold selection procedure, it is unusual to consider thresholds which lie below an obvious mode. Many threshold selection procedures are simply not set up to handle data from non-monotonically decreasing densities. So we explore a different choice for the candidate threshold grids here to ensure that our choices made in the main paper do not unfairly favour the EQD method.
Given that our aim is to provide a method which requires no user input, we consider candidate thresholds across the range of the sample data and allow the method to make the threshold selection whether the conditions are suitable or not. In the main paper and the supplementary material, outside of Gaussian case where we specifically restricted our choice of candidate thresholds such that $u>q_{50}$ (with $q_{50}$ being the sample median), and Case 0, where the optimal threshold is at the lowest sample quantile, we have only explored cases where candidate thresholds span the range of the dataset. Here, we reanalyse two cases; Case 1 and Case 4, now with candidate threshold grids which lie above the mode of the distributions.
To avoid sensitivities from different mode estimators clouding the results, we use the true mode to help define our range/set of candidate thresholds. In Case 1, the true mode is trivial to find and it is the optimal threshold. For Case 4, the mode needs to be found numerically, using the expression for the density function. Once the mode is known, we define the candidate threshold set as sample quantiles at levels $0\%,\ldots,95\%$ of the data which lies above the mode.
First, we consider the effect of the range $[u_{1},u_{k}]$ of the candidate thresholds on threshold selection performance. Table S.20 provides the RMSE, bias and variance of threshold estimation for samples from both of these cases, using the EQD, Wadsworth and Northrop methods each applied with candidate thresholds across the distribution, as given in Section 6.1 of the main paper, and only above the true mode. In terms of RMSE, the EQD method outperforms the other two methods regardless of how the candidate threshold grid is chosen. For the “Above mode” grid, we have rather different performances for the Wadsworth and Northrop methods relative to the “Original” candidate grid. In Case 1, these two methods perform very similarly, with RMSEs approximately 15 times larger than the EQD. In Case 4, the Wadsworth method fails completely with the “Above mode” grid due to the grid being too fine relative to the sample size. For the Northrop method, the RMSE of threshold estimation is increased with the “Above mode” grid and the differential between the Northrop and EQD becomes wider with a RMSE that is 1.15 times larger the EQD method.
RMSE Bias Variance RMSE Bias Variance RMSE Bias Variance
Case 1
EQD Wadsworth Northrop
Original 0.048 0.034 0.001 0.349 0.111 0.110 0.536 0.276 0.212
Above mode 0.038 0.018 0.001 0.577 0.236 0.277 0.597 0.348 0.235
Case 4
Original 0.526 $-0.515$ 0.012 0.627 $-0.407$ 0.230 0.543 $-0.222$ 0.246
Above mode 0.514 $-0.505$ 0.009 NA NA NA 0.589 $-0.127$ 0.331
Table S.20: RMSE, bias and variance of threshold choice for Case 1 and Case 4 samples for EQD, Wadsworth and Northrop methods. The methods are evaluated for two sets of candidate threshold sets: Original, using candidate threshold across the whole sample; Above mode, using candidate thresholds which only lie above the mode.
For the EQD method, using only candidate thresholds above the true mode results in a smaller RMSE (and bias and variance) for both Cases 1 and 4, presumably as we are now using a finer grid of candidate thresholds so candidate thresholds lie closer to the optimal value. In contrast, the performance of the Wadsworth and Northrop methods is worsened when we restrict the candidate thresholds to be above the mode, primarily due to an increase in variance.
We also consider the effect of the range of the candidate thresholds on quantile estimation for the Gaussian case. Previously, for the Gaussian case, we had reported results only using candidate threshold above the sample median of the data, i.e., essentially the mode. Here, we assess the effect of allowing candidate thresholds to range across the whole sample, i.e., the sample quantiles at levels $0\%,5\%,\ldots,95\%$ of the whole data sample.
The results, in terms of RMSE for three different quantiles, are presented in Table S.21 for the two different candidate threshold grids. Firstly, we find that the Wadsworth method completely fails with this extended range of candidate thresholds, so this method is omitted from the table. For both the EQD and Northrop methods, we have almost identical values regardless of the set of candidate thresholds we use, with the RMSE always smallest for the EQD.
EQD Northrop
$j$ Original Whole data Original Whole data
0 0.214 0.214 0.225 0.225
1 0.430 0.431 0.461 0.460
2 0.703 0.707 0.765 0.763
Table S.21: RMSEs of estimated $(1-p_{j,n})$-quantiles where $p_{j,n}=1/(10^{j}n)$, for $j=0,1,2$, for the Gaussian case when applying methods with candidate thresholds across the range of the data.
The results here indicate that the choice of candidate grid taken in the main paper provides no unfair advantage to the EQD method, in fact it performs even better if the choice of candidate threshold grid is tuned by exploiting knowledge of the value of the mode of the distribution.
@@
Tables S.22 and S.23 expanding on the results given in Tables 4 and 6 of the main paper, providing coverage probabilities and average CI width ratios using Algorithms 1 and 2 for 500 samples derived from Case 4 and the Gaussian distribution. These results allow us to conclude that the additional threshold uncertainty, captured by Algorithm 2, leads to a significant improvement in the calibration of CIs and that it is vital to include this uncertainty in any inference for extreme quantiles.
@@
Table S.22 shows that for all confidence levels and exceedance probabilities, Algorithm 1, which includes uncertainty only from the GPD parameter estimation, substantially under-estimates the uncertainty in quantile estimates. This leads to CIs which do not cover the true quantiles to the nominal levels of confidence. In contrast, the inclusion of the additional threshold uncertainty in Algorithm 2 leads to significant increases in the coverage of true quantiles with coverage probabilities at all quantile levels lying very close to the nominal confidence level, especially at the 95% level. From a practical perspective, it is reassuring that this improvement in coverage is achieved with only 40-68% average increases in the width of the CIs.
@@
Table S.23 shows that, for Gaussian variables across all quantiles, Algorithm 1 and 2 are less successful in the coverage of the true quantiles than in Case 4. Specifically, there is a significant under-estimation of the estimated uncertainty necessary to provide coverage probabilities near to the nominal confidence level, and the actual coverage is decreasing with the level of extrapolation required. This is not too surprising as it is well-established that Gaussian variables exhibit quite slow convergence to an extreme value limit. However, what Table S.23 shows is that the additional threshold uncertainty in Algorithm 2 leads to a substantial improvement in actual coverage across all quantiles and for all nominal confidence levels. This improvement is achieved with the CI widths on average being extended by 45-73%. In particular, for the $p=1/n,1/3n,1/5n$, which are typical levels of extrapolation from a sample in practical contexts, Algorithm 2 achieves a workable performance, with coverage reasonably close to the nominal level.